Can you help me with this problem please? 2u+v=9
u/2 -1/2=-v
I have to solve it using elimination but every time I try to check my solution it comes out incorrect. Can you help me please? I would really appreciate it! Thank you!

Answer by edjones(8007)

2u+v=9
.
u/2 -1/2=-v
u-1=-2v multiply each side by 2.
u+2v=1
-2u-4v=-2 multiply each side by -2.
.
2u+v=9
-2u-4v=-2
-----------add
-3v=7
.
v=-7/3
u=17/3
.
Ed